Solve the system of linear equations by the substitution method.
5x + 4y = 14
15x - 2y = 7

Answer by Alan3354(69443)

Solve the system of linear equations by the substitution method.
5x + 4y = 14
15x - 2y = 7
-------------
-2y = 7-15x (from 2nd eqn)
4y = 30x-14
Sub into 1st eqn
5x + (30x-14) = 14
35x - 14 = 14
35x = 28
x = 28/35 = 4/5
------------
Sub into either, I'll use eqn 1
5*(4/5) + 4y = 14
4 + 4y = 14
4y = 10
y = 10/4 = 5/2